OpenStudy (anonymous):

If , the value of D=m over v when D = 5.1 and v = 0.3 is ------

OpenStudy (jackellyn):

D=m/v 5.1 = m/0.3 It seems you want to solve for m Multiply both sides by 0.3, it will cancel it on the right side and give you the value of m once you multiply it by the left side. 5.1(0.3)=m

OpenStudy (jackellyn):

@waterineyes, I find it interesting that you gave me a lot of grief about giving the answer to the person asking the question, and that you did the same thing. Just pointing it out, if you're going to give advice, you should also follow it. Thanks! :)

OpenStudy (anonymous):

What?? What are you saying I am not getting?? Where I gave the answer??? Ha ha ha... Okay Sorry..